# Continuous Speech Synthesis using per-token Latent Diffusion

### v1

@@ Abstract
The success of autoregressive transformer models with discrete tokens has inspired quantization-based approaches for continuous modalities, though these often limit reconstruction quality. We therefore introduce SALAD, a per-token latent diffusion model for zero-shot text-to-speech, that operates on continuous representations. SALAD builds upon the recently proposed expressive diffusion head for image generation, and extends it to generate variable-length outputs. Our approach utilizes semantic tokens for providing contextual information and determining the stopping condition. We suggest three continuous variants for our method, extending popular discrete speech synthesis techniques. Additionally, we implement discrete baselines for each variant and conduct a comparative analysis of discrete versus continuous speech modeling techniques. Our results demonstrate that both continuous and discrete approaches are highly competent, and that SALAD achieves a superior intelligibility score while obtaining speech quality and speaker similarity on par with the ground-truth audio.
1 Introduction
Autoregressive (AR) modeling is often correlated with discrete representations, probably due to the remarkable success of Large Language Models (LLMs), which operate on a discrete modality. (Vaswani et al., 2017; Radford et al., 2019). Inspired by the success of LLMs, continuous modalities, such as audio and images, are quantized to be modeled discretely. In image generation, quantization is often achieved by discrete autoencoders (Van Den Oord et al., 2017), which are later optimized with adversarial losses to improve fidelity (Esser et al., 2021). Works that focus on audio generation usually employ Residual Vector Quantization (RVQ) (Zeghidour et al., 2021), a process that iteratively refines the approximation by quantizing the residual. The resulting discrete codes are used for discrete AR modeling (Esser et al., 2021; Wang et al., 2023; Copet et al., 2024).
Discrete modeling over continuous domains requires quantization, which degrades the reconstruction quality and upper-bounds the fidelity. Using multiple RVQ quantizers enhances the fidelity, but the fine RVQ codes might quantize noise, which can be detrimental for discrete modeling methods. Discrete autoencoders may also suffer from low codebook utilization (Mentzer et al., 2023), and multimodal models that work on discrete representation suffer from stability issues (Team, 2024). We therefore suspect that quantizing inherently continuous modalities may be sub-optimal, and focus on continuous alternatives instead.
Predicting continuous distributions with regression losses such as L1 or L2, induce a unimodal distribution, an unrealistic assumption for most generative tasks. We hypothesize that multimodal distributions, which enable multiple local maxima, can represent more complex patterns and is crucial for generative one-to-many tasks. Recent works in image generation have explored approaches to modeling continuous distributions. GIVT (Tschannen et al., 2023) represents the continuous distribution using a Gaussian Mixture Model, while AR-Diffusion (Li et al., 2024) suggests a per-token diffusion head to model the continuous frame distributions.
@@ 1 Introduction
We suggest SALAD (Speech synthesis with Autoregressive LAtent Diffusion), a per-token latent diffusion model for zero-shot speech synthesis over continuous representations, inspired by the per-token diffusion head suggested by Li et al. (2024). We enable the generation of variable length outputs, addressing a challenge that is absent in image generation methods, where the number of tokens to generate is fixed. We utilize semantic tokens (Kharitonov et al., 2023; Borsos et al., 2023a) – quantized embeddings of a self-supervised model – for contextual information and to define the generation-stopping condition. SALAD does not rely on text-audio alignment, which makes it easier to leverage large data sources, and can synthesize audio based on a target speaker using a three-second speaker prompt. We propose three variants for SALAD:
  1. 1.
    T2A (Text2Acoustic): predicts acoustic features directly from text, using semantic tokens as an auxilary task.
  2. 2.
    S2A-AR (Semantic2Acoustic Autoregressive): predicts acoustic features from semantic tokens by next-token prediction.
  3. 3.
    S2A-NAR (Semantic2Acoustic Non-Autoregressive): predicts acoustic features from semantic tokens using a MaskGIT (Chang et al., 2022) schedule.
For each of our continuous variants, we train a comparable model that operates on discrete representations, replacing the diffusion head with RVQ discrete prediction heads (Figure 1). Our discrete T2A model is the first to predict semantic and acoustic tokens in parallel directly from text. Our SoundStorm discrete baseline employs the random unmasking method, which we demonstrate to outperform the confidence-based unmasking. We evaluate all models on speech quality, intelligibility, and speaker similarity. The results suggest that SALAD’s T2A model achieves the highest intelligibility score, while having speech quality and similarity scores on-par with the ground-truth audio, as measured in subjective listening tests  [^1] Samples are available at https://s3.us-south.objectstorage.softlayer.net/zk-wav-data/Webpages/ICLR2025PerTokenLatentDiffusion/index.html. Our contributions can be summarized as follows:
  • Propose SALAD, a zero-shot speech synthesis system that uses per-token latent diffusion.
  • Extend popular discrete speech synthesis methods to continuous representations.
  • Suggest a discrete text-to-acoustic model and improve SoundStorm’s unmasking method.
  • Compare discrete and continuous modeling techniques in a controlled environment.
2 Related Work
Zero-Shot TTS
@@ 2 Related Work
Inspired by the success of in-context learning, there has been great interest in text-to-speech (TTS) systems that can generalize to unseen speakers during inference. This task is commonly known as zero-shot TTS, and provides many benefits due to its flexibility and increased quality (Wang et al., 2023). Zero-shot TTS systems, including SALAD, typically formulate the problem as a language modeling task, containing text and audio tokens. Such methods make use of a speaker prompt – a short recording from the target speaker – and synthesize the text according to the prompt (Le et al., 2024; Shen et al., 2023; Łajszczak et al., 2024; Peng et al., 2024).
Semantic Tokens
Quantized embeddings of self-supervised audio models (Hsu et al., 2021; Baevski et al., 2020; Chung et al., 2021), commonly known as semantic tokens, have been shown to capture phonetic and prosodic information, and show benefits when modeling long-range dependencies in various speech processing tasks. These tokens became popular as intermediate representations for speech synthesis (Kharitonov et al., 2023; Huang et al., 2023; Borsos et al., 2023a), unconditional audio generation (Borsos et al., 2023b), and for text-audio multimodal tasks (Rubenstein et al., 2023). SALAD predicts semantic tokens as an auxilary task to obtain contextual information and determine the stopping condition.
RVQ codes prediction
Various audio generation methods have designed unique ways to predict the RVQ codes matrix, each with their advantages and limitations. AudioLM (Borsos et al., 2023b) flattens the codes matrix into a long sequence, which greatly increases the token count in transformer models. Most followup works avoid flattening and embed all RVQ residual layers into a single token. Vall-E (Wang et al., 2023) generates the initial coarse code vector with an AR model and then uses an non-autoregressive (NAR) model to predict the rest of the codebooks. AudioGen (Kreuk et al., 2022) uses a single AR model to predict all codes of each timestep in parallel. MusicGen (Copet et al., 2024) extends AudioGen by introducing a delay pattern, ensuring each code is predicted based on its coarser RVQ layers, leading to a better approximate factorization. SoundStorm (Borsos et al., 2023a) employs the fast NAR decoding algorithm by MaskGIT (Chang et al., 2022) to generate acoustic tokens based on semantic tokens. NaturalSpeech3 (Ju et al., 2024) trains a factorized codec, which disentangles speech characteristics into discrete factors and predicts each factor using a MaskGIT procedure. As opposed to all above methods, SALAD directly predicts a continuous latent space, thus avoiding the need to predict multiple residuals codes.
Continuous models
@@
When learning a continuous distribution, recent works typically use diffusion models, which were developed to sample from complex continuous probability distributions, inspired by non-equilibrium thermodynamics (Ho et al., 2020). Several works attempt to synthesize speech using a diffusion process, which has the challenge of generating variable length outputs (Kong et al., 2020; Chen et al., 2020; Popov et al., 2021). For that end, most diffusion-based works rely on a duration predictor that predicts the audio length in advance, which might be inferior to determining the length on-the-fly during synthesis (Shen et al., 2023; Le et al., 2024). MELLE (Meng et al., 2024) predicts Mel spectrograms autoregressively using a Gaussian sampling module, and parameterizes the next frame using a Gaussian distribution, which restricts it to learn only unimodal distributions. MELLE relies on an additional binary classifier that indicates when to stop, which is a highly imbalanced classification problem. In contrast, SALAD operates on VAE latent tokens, which allows sampling diverse inputs while training, and uses a diffusion head, capable of modeling multimodal distributions. SALAD relies on semantic-tokens to determine the stopping condition, a more balanced representation which also provides contextual information.
3 Method
3.1 Background
Definitions
We denote the raw audio sequence as $\bm{a}=(a_{1},...,a_{m})$ where $a_{i}\in[-1,1]$ with sampling rate $f_{S}$. The text is $\bm{y}=(y_{1},..,y_{k})$ where $y_{i}\in\mathcal{A}$, and $\mathcal{A}$ is the text vocabulary. We obtain compressed audio representations using a variational autoencoder (VAE), trained with adversarial losses to obtain high-fidelity reconstructions. The VAE’s encoder $\mathcal{E}$ predicts a sequence of means and variances of normal distribution: $(\mu_{1},...,\mu_{n}),(\sigma^{2}_{1},...,\sigma^{2}_{n})=\mathcal{E}(\bm{a})$ where $\sigma_{i},\mu_{i}\in\mathbb{R}^{d}$ and $d$ is the VAE bottleneck dimension. The VAE downsamples the sequence with a stride $r$. We sample $x_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$ and denote $\bm{x}=(x_{1},..,x_{n})$ as the continuous acoustic tokens. The VAE’s decoder $\mathcal{D}$ is used for reconstruction $\hat{a}_{1},...,\hat{a}_{m}=\mathcal{D}(x_{1},..,x_{n}).$ We also extract semantic tokens and denote them by $\bm{w}=(w_{1},..,w_{m})$, which have the same downsampling stride as the VAE. Our goal is to predict the audio based on the desired text and the speaker prompt. Denoting the speaker prompt latent features as $\bm{s}=s_{1},...,s_{p}$, our training objective can be formulated by: $p(\bm{x}|\bm{y},\bm{s})$.
Diffusion Process
A diffusion process starts from a continuous signal, and gradually destroys it using a forward noise process. Our method performs latent diffusion, and attempts to predict the VAE latent vectors $x_{1},...,x_{n}$. Given noising coefficients $\beta_{0},...,\beta_{T}$ and some continuous vector $x$, we define $x^{0}=x$ and $\epsilon\sim\mathcal{N}(0,I)$; the Markov structure is $x^{t}=\sqrt{1-\beta_{t}}x^{t-1}+\sqrt{\beta_{t}}\epsilon$. This iterative denoising process can be simplified. By defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod^{t}_{i=1}\alpha_{i}$, we get that $x^{t}=\sqrt{\bar{\alpha_{t}}}x+\sqrt{1-\bar{\alpha_{t}}}\epsilon$. The diffusion process is often defined such that $\bar{\alpha}_{T}\to 0$ and $x^{T}$ distributes closely to the standard normal distribution. Diffusion models $\epsilon_{\theta}$ are trained to perform the reverse diffusion process, which denoises the corrupted signal by predicting the added noise. Their denoising loss is defined as $\mathcal{L}(x)=\mathbb{E}_{\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(t,x^{t})\|^{2}\right]$. Most diffusion models operate on a sequence $x_{1},..,x_{n}$ and attempt to denoise all tokens in parallel using $\epsilon_{\theta}(t,x_{1}^{t},...,x_{n}^{t})$.
@@ 3.1 Background
Per-Token Diffusion Head
@@
Li et al. (2024) proposed an MLP diffusion head for image generation. Unlike standard diffusion models, the diffusion head denoises each token independently, which gives additional flexibility when defining the conditioning information (e.g., predicting on previously predicted tokens). We rely on a transformer model $\Theta$ that extracts contextual per-token conditioning vectors $z_{1},..,z_{n}$ based on the input features and optional context vectors that we denote by $C$
$\bm{z}=z_{1},...,z_{n}=\Theta(C,x_{1},...,x_{n})$
@@
The diffusion head (noise estimator) $\epsilon_{\theta}$ takes a contextual conditioning vector $z$ and attempts to model the continuous distribution $p(x|z)$. Given a target token $x$, we follow a similar diffusion process but condition the prediction on $z$. The loss is
$\mathcal{L}(x,z)=\mathbb{E}_{\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(x^{t},t,z)\|^{2}\right]$ (1)
@@
During training, we sample $t\sim[T],\epsilon\sim\mathcal{N}(0,I)$ for each token $x$, obtain the noisy targets $x^{t}$, and minimize $\mathcal{L}(x,z)$ (Figure 2(a)). This denoising network is trained jointly with the transformer $\Theta$, and the gradient with respect to $z$ is propagated to the transformer. We can sample $K$ different values of $t,\epsilon$ for a given context vector and target $z,x$, with the additional complexity of just the MLP head rather than the entire model. During inference, we sample a continuous vector by sampling a Gaussian vector $x^{T}\sim\mathcal{N}(0,I)$ and reversing the diffusion process (see Figure 2(b)):
@@
$x^{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x^{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x^{t},t,z)\right)+\sqrt{\beta_{t}}\epsilon$ (2)
3.2 SALAD: Speech synthesis using Autoregressive LAtent Diffusion

### v2

# Speech Synthesis From Continuous Features Using Per-Token Latent Diffusion

@@ Abstract
We present SALAD, a zero-shot text-to-speech (TTS) autoregressive model operating over continuous speech representations. SALAD utilizes a per-token diffusion process to refine and predict continuous representations for the next time step. We compare our approach against a discrete variant of SALAD as well as publicly available zero-shot TTS systems, and conduct a comprehensive analysis of discrete versus continuous modeling techniques. Our results show that SALAD achieves superior intelligibility while matching the speech quality and speaker similarity of ground-truth audio.
@@
I Introduction
Discrete autoregressive (AR) models have become the leading approach in natural language processing. Their remarkable success has motivated researchers to extend these models to image and speech domains. Modeling continuous domains with discrete approaches typically involves quantization. In image generation, quantization is often realized using discrete autoencoders [VQVAE], which are further optimized with adversarial losses [VQGAN, DALLE, Make-a-scene]. In audio generation, techniques like Residual Vector Quantization (RVQ) [Soundstream, Encodec] iteratively quantize the residual, producing multiple discrete codes per frame that can be modeled autoregressively [VQGAN, VALLE, Musicgen].
While presenting impressive results, discrete approaches have several limitations. First, modeling multiple dependent streams simultaneously is inherently complex and often requires additional strategies like two-stage modeling [VALLE, VallE2, VallEX] or token interleaving [Musicgen]. Second, similarity among consecutive codec codes has been shown to cause robustness issues, leading to continuous stretches of silence or persistent noise [VallE2]. Third, quantization methods inherently impose a fidelity upper-bound, constrained by continuous representations such as mel-spectrogram features [puvvada2024discrete]. This raises the question: Can we optimize an AR model to directly generate high-quality speech over continuous features?
Recently, several studies have explored language modeling with continuous-valued tokens. In image generation, GIVT [GIVT] proposes representing the continuous distribution of images obtained by a VAE using a Gaussian Mixture Model (GMM), while AR-Diffusion [ARDiffusion] suggests a per-token image diffusion head to model similar representations. For speech generation, [MELLE] presented an AR modeling approach over mel spectrograms, incorporating a Gaussian sampling module followed by a post-net. Concurrently with our work, [lin2025continuous] propose a similar modeling approach to GIVT but for spoken data.
Inspired by AR-Diffusion [ARDiffusion], we propose SALAD (Speech synthesis with Autoregressive LAtent Diffusion), an autoregressive per-token latent diffusion model for zero-shot speech synthesis over continuous features. By operating directly in continuous space, SALAD eliminates the need for signal quantization, thus enabling higher-fidelity speech generation. As an autoregressive model, SALAD also naturally supports variable-length outputs, addressing a fundamental challenge not encountered in fixed-length image generation methods. We use semantic tokens [Hubert, w2v, w2vbert] which are not used for signal synthesis but rather to enhance robustness and define the generation-stopping condition.
We evaluate SALAD against the discrete representation alternative employing RVQ quantization, specifically following the delay prediction pattern introduced by [Musicgen] (see Figure 1). Additionally, we compare SALAD with two leading publicly available zero-shot TTS systems, XTTS [XTTS] and VoiceCraft [peng2024voicecraft]. Our experimental results demonstrate that SALAD achieves superior intelligibility and maintains speech quality and speaker similarity comparable to ground-truth audio, as confirmed by both objective metrics and subjective listening tests  [^1] Samples available at: https://s3.us-south.objectstorage.softlayer.net/zk-wav-data/Webpages/SynthesisPerTokenLatentDiffusion/index.html.
Our contributions:  (i) We introduce SALAD, the first autoregressive per-token latent diffusion model for zero-shot speech synthesis directly over continuous acoustic representations, eliminating the need for discrete quantization. (ii) We empirically demonstrate that continuous latent diffusion modeling can surpass discrete modeling approaches in intelligibility and match them in quality and speaker similarity. (iii) Through comprehensive experiments and ablation studies, we analyze the advantages and trade-offs of continuous versus discrete speech modeling, providing practical insights into the design of zero-shot speech generation models.
II Related Work
@@
Zero-Shot TTS.  Inspired by the success of in-context learning, there has been significant interest in zero-shot TTS systems that generalize to unseen speakers during inference, offering flexibility and improved quality [VALLE]. Zero-shot TTS methods typically formulate the task as language modeling over text and audio tokens, leveraging short speaker prompts to synthesize speech aligned with the target speaker’s voice characteristics [Voicebox, NaturalSpeech2, BaseTTS, peng2024voicecraft]. Our method follows this paradigm but operates directly over continuous speech representations.
Semantic Tokens.  Quantized embeddings from self-supervised audio models, termed semantic tokens [Hubert, w2v, w2vbert], capture phonetic and prosodic information beneficial for speech synthesis [SPEAR, MakeAVoice, Soundstorm], unconditional audio generation [AudioLM], and multimodal text-audio tasks [AudioPalm]. SALAD utilizes semantic tokens as an auxiliary representation defining a precise stopping condition for continuous acoustic modeling.
RVQ Codes Prediction.  Discrete audio generation methods commonly utilize Residual Vector Quantization (RVQ), involving multiple quantization layers [Soundstream, Encodec]. Methods such as AudioLM [AudioLM] flatten RVQ codes into sequences, whereas others, like Vall-E [VALLE], employ two-stage prediction approaches or token interleaving strategies [Musicgen]. Differently, SALAD  directly predicts a continuous latent space, thus avoiding the need to predict multiple residuals codes.
@@
Continuous Models.  When learning a continuous distribution, recent works typically use diffusion models, which were developed to sample from complex continuous probability distributions, inspired by non-equilibrium thermodynamics [DDPM]. Several works attempt to synthesize speech using a diffusion process, which has the challenge of generating variable length outputs [Diffwave, Wavegrad, GradTTS]. For that end, most diffusion-based works rely on a duration predictor that predicts the audio length in advance, which might be inferior to determining the length on-the-fly during synthesis [NaturalSpeech2, Voicebox]. MELLE [MELLE] predicts Mel spectrograms autoregressively using a Gaussian sampling module, and parameterizes the next frame using a Gaussian distribution, which restricts it to learn only unimodal distributions. MELLE relies on an additional binary classifier that indicates when to stop, which is a highly imbalanced classification problem. In contrast, SALAD operates on VAE latent tokens, which allows sampling diverse inputs while training, and uses a diffusion head, capable of modeling multimodal distributions. SALAD relies on semantic-tokens to determine the stopping condition, a more balanced representation which also provides contextual information.
@@
III Background
Definitions.  We denote the raw audio sequence as $\boldsymbol{a}=(a_{1},...,a_{m})$ where $a_{i}\in[-1,1]$ with sampling rate $f_{S}$. The text is $\boldsymbol{y}=(y_{1},..,y_{k})$ where $y_{i}\in\mathcal{A}$, and $\mathcal{A}$ is the text vocabulary. We obtain compressed audio representations using a variational autoencoder (VAE), trained with adversarial losses to obtain high-fidelity reconstructions. The VAE’s encoder $\mathcal{E}$ predicts a sequence of means and variances of normal distribution: $(\mu_{1},...,\mu_{n}),(\sigma^{2}_{1},...,\sigma^{2}_{n})=\mathcal{E}(\boldsymbol{a})$ where $\sigma_{i},\mu_{i}\in\mathbb{R}^{d}$ and $d$ is the VAE bottleneck dimension. The VAE downsamples the sequence with a stride $r$. We sample $x_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$ and denote $\boldsymbol{x}=(x_{1},..,x_{n})$ as the continuous acoustic tokens. Hereafter, we refer to these continuous VAE latents interchangeably as acoustic tokens or continuous acoustic representations, following prior usage in speech-diffusion literature[MELLE]. The VAE’s decoder $\mathcal{D}$ is used for reconstruction $\hat{a}_{1},...,\hat{a}_{m}=\mathcal{D}(x_{1},..,x_{n}).$ We also extract semantic tokens and denote them by $\boldsymbol{w}=(w_{1},..,w_{m})$, which have the same downsampling stride as the VAE. Our goal is to predict the audio based on the desired text and the speaker prompt. Denoting the speaker prompt latent features as $\boldsymbol{s}=s_{1},...,s_{p}$, our training objective can be formulated by: $p(\boldsymbol{x}|\boldsymbol{y},\boldsymbol{s})$.
Diffusion Process.  A diffusion process starts from a continuous signal, and gradually destroys it using a forward noise process. Our method performs latent diffusion, and attempts to predict the VAE latent vectors $x_{1},...,x_{n}$. Given noising coefficients $\beta_{0},...,\beta_{T}$ and some continuous vector $x$, we define $x^{0}=x$ and $\epsilon\sim\mathcal{N}(0,I)$; the Markov structure is $x^{t}=\sqrt{1-\beta_{t}}x^{t-1}+\sqrt{\beta_{t}}\epsilon$. This iterative denoising process can be simplified. By defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod^{t}_{i=1}\alpha_{i}$, we get that $x^{t}=\sqrt{\bar{\alpha_{t}}}x+\sqrt{1-\bar{\alpha_{t}}}\epsilon$. The diffusion process is often defined such that $\bar{\alpha}_{T}\to 0$ and $x^{T}$ distributes closely to the standard normal distribution. Diffusion models $\epsilon_{\theta}$ are trained to perform the reverse diffusion process, which denoises the corrupted signal by predicting the added noise. Their denoising loss is defined as $\mathcal{L}(x)=\mathbb{E}_{\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(t,x^{t})\|^{2}\right]$. Most diffusion models operate on a sequence $x_{1},..,x_{n}$ and attempt to denoise all tokens in parallel using $\epsilon_{\theta}(t,x_{1}^{t},...,x_{n}^{t})$.
@@ III Background
Per-Token Diffusion Head.  [ARDiffusion] proposed an MLP diffusion head for image generation. Unlike standard diffusion models, the diffusion head denoises each token independently, which gives additional flexibility when defining the conditioning information (e.g., predicting on previously predicted tokens). The authors rely on a transformer model $\Theta$ that extracts contextual per-token conditioning vectors $z_{1},..,z_{n}$ based on the input features and optional context vectors that we denote by $C$
$\boldsymbol{z}=z_{1},...,z_{n}=\Theta(C,x_{1},...,x_{n}).$
@@
The diffusion head (noise estimator) $\epsilon_{\theta}$ takes a contextual conditioning vector $z$ and attempts to model the continuous distribution $p(x|z)$. Given a target token $x$, a diffusion process is being applied conditioned on $z$. The loss is:
$\mathcal{L}(x,z)=\mathbb{E}_{\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(x^{t},t,z)\|^{2}\right].$ (1)
@@
During training, $t\sim[T],\epsilon\sim\mathcal{N}(0,I)$ is sampled for each token $x$, resulting in noisy targets $x^{t}$, and $\mathcal{L}(x,z)$ is minimized. See Figure 2(a). The denoising network is trained jointly with the transformer $\Theta$, and the gradient with respect to $z$ is propagated to the transformer. $K$ different values of $t,\epsilon$ may be sampled for a given context vector and target $z,x$, with the additional complexity of just the MLP head rather than the entire model. During inference, a continuous vector is sampled from $x^{T}\sim\mathcal{N}(0,I)$ and the final result is obtained from the reverse diffusion process (see Figure 2(b)):
@@
$x^{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x^{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x^{t},t,z)\right)+\sqrt{\beta_{t}}\epsilon.$ (2)
Figure 3: System overview. Conditioned on text $t$ and a speaker prompt $s$, SALAD emits semantic tokens $w$ and acoustic latents $x$, which a VAE decoder converts to waveform.
IV SALAD